1.5 and Q.1.6 are based on the following information about an economy that
produces the following combinations of two goods: bats and balls.
2018 (Base year) 2019
Quantity Price (R) Quantity Price (R)
Bats 205 8 210 9
Balls 310 4 320 5
Q.1.5 The value of the nominal GDP in 2019 is: (2)
(1) R1 530.
(2) R1 045.
(3) R2 880.
(4) R3 490.
Q.1.6 The value of the real GDP in 2019 is: (2)
(1) R1 880.
(2) R2 620.
(3) R2 960.
(4) R3 395.

Q.1.7 If the marginal propensity to consume
1
Expert's answer
2020-10-02T06:42:16-0400
"Solution"

Q. 1.5 Nominal GDP in 2019:


"=current\ year\ Price\ \cdot Quantity\\
=(210 \cdot 9)+(320 \cdot 5) \implies 1890+1600=R3490"

The solution is option (4)


Q.1.6 The value of the real GDP in 2019:


"=Base\ year\ Price\ \cdot Current\ Quantity\\
=(210 \cdot 8)+(320 \cdot 4) \implies 1680+1280=R2960"

The solution is option (3)


Q 1.7 marginal propensity to consume increases

This means change in consumption is more than change in income so consumption curve slope become steeper